# Spike Encoding for Environmental Sound: A Comparative Benchmark

### v3

@@ 1 Introduction
Spiking Neural Networks (SNNs) have emerged as a promising paradigm in neuromorphic computing due to their energy efficiency and biologically inspired processing mechanisms [1]. Unlike traditional artificial neural networks (ANNs), which rely on continuous-valued activations [2], SNNs process information through discrete spike events, mimicking the behavior of biological neurons [3]. To harness SNNs for real-world applications, sensor data need to be represented as spike trains. While neuromorphic sensors are specifically designed to generate spike-based outputs directly [4], conventional sensors typically require an additional encoding step. In this context, various spike encoding methods have been proposed, each with distinct characteristics influencing the efficiency and accuracy of SNN-based models [5]. The selection of an appropriate encoding scheme is crucial, as it directly affects the network’s capability to extract meaningful temporal patterns from input signals [6].
Although spike encoding techniques have been extensively studied in the context of speech processing [7], their application to environmental sound processing remains underexplored. Environmental sounds, such as traffic noise, alarms, and natural ambient sounds, exhibit complex spectral-temporal structures, often containing overlapped acoustic events and substantial background noise [8]. Unlike speech, which follows structured linguistic patterns, environmental sounds are highly unstructured and vary significantly in duration and frequency content. These unique challenges require a deeper investigation into the suitability of different spike encoding schemes for environmental sound processing.
State-of-the-art environmental sound classification methods predominantly employ convolutional neural networks (CNNs) or other machine learning models applied to spectrogram representations, which effectively capture the inherent time-frequency characteristics of audio signals [9, 10, 11, 12]. However, to the best of our knowledge, no state-of-the-art solution has yet encoded environmental sound datasets using spike-based methods for SNNs. This gap presents a promising opportunity to investigate neuromorphic-inspired approaches that may offer good temporal resolution and enhanced energy efficiency.
@@ 2.1 Datasets
ESC-10: A subset of ESC-50 containing 400 recordings, divided into 10 classes with 40 clips per class. Each clip has a duration of 5 s. The dataset provides a balanced representation of environmental events, including dog barking, rain, sea waves, crying baby, clock ticking, sneezing, helicopter, chainsaw, crowing rooster, and fire crackling. We used all original 5 s recordings without modification [13].
UrbanSound8K: The original dataset contains 8,732 recordings distributed across 10 classes (air conditioner, car horn, children playing, dog bark, drilling, engine idling, gun shot, jackhammer, siren, and street music), each clip lasting up to 4 s. To standardize audio length and ensure a fair comparison in the classification analysis, we retained only clips of exactly 4 s. Having the same lengths avoids padding or other manipulations that could bias the results. We also maintained class balance across folds. However, due to an insufficient number of valid 4 s samples, the classes drilling and car horn were removed, resulting in a filtered subset containing 8 classes and 3,768 clips [14].
TAU-3Class: Derived from the DCASE 2020 Acoustic Scene Classification task, we used recordings from device A grouped into three categories: indoor, outdoor, and transportation. Each original recording is 10 s. To match lengths with the other datasets, we extracted the central 5 s segment, ensuring retention of the most relevant acoustic information. We subsampled the official development split to balance the classes and simplify the analysis, yielding 1,500 training and 300 test clips (1,800 total) [15].
@@ 2.2 Signal Reconstruction and Multi-band Analysis
The Mel-spectrogram was chosen for its computational efficiency and demonstrated effectiveness [7]. While prior work has employed Gammatone or Butterworth filters to compute cochleagrams [16, 17], our tests indicated higher computational costs for high-frequency content; the Mel-spectrogram provided a more efficient solution for environmental sound analysis.
@@
Reconstruction quality was assessed using two metrics computed per frequency band—ERRdB (error in decibels) and SNR (signal-to-noise ratio). In addition, ERRdB was computed per sound class for finer-grained analysis. Signal reconstruction was performed using the decoding methods described in [6].
@@ 2.3 Spike Encoding Methods
Spiking neural networks (SNNs) require transforming continuous signals into discrete spike trains, a process known as spike encoding. The choice of encoding method strongly influences downstream performance. In this study, three techniques—Moving Window (MW), Step Forward (SF), and Threshold Adaptive Encoding (TAE)—were selected based on demonstrated effectiveness in prior work [17, 6, 18], yet remain underexplored for environmental sound classification, motivating their evaluation here.
@@
Moving Window (MW): Partitions the signal into overlapping windows and emits spikes based on intra-window signal variation, capturing local temporal patterns [17].
Step Forward (SF): Generates spikes when the difference between consecutive samples exceeds a predefined threshold, preserving salient changes while suppressing minor fluctuations [17].
Threshold Adaptive Encoding (TAE): Adapts the spike-generation threshold to signal characteristics, improving robustness to amplitude variability and sensitivity to significant changes [18].
@@ 2.4 SNN Classification
To evaluate classification accuracy, we implemented a spiking neural network (SNN) with four fully connected layers and Leaky Integrate-and-Fire (LIF) neurons using snntorch (v0.9.1) [2]. The architecture comprises an input layer matching the Mel-spectrogram channels, three hidden layers with 128 neurons each, and an output layer matching the number of classes. Training was performed with a batch size of 32, a learning rate of 0.01, and 100 epochs.
@@ 3.3 Classification
  • Note: Baseline references values taken from: ESC-10 [13]; UrbanSound8K [14]; TAU-3Class (Task B) [19].
Our results highlight TAE as the most effective spike encoding method for both reconstruction and classification of environmental sounds, supporting its robust generalization and efficiency. This corroborates prior studies on adaptive thresholding for sequential data [18] and suggests future work should explore further architectural improvements for challenging sound recognition tasks.
@@ 4 Discussion
Our multi-faceted analysis reveals that Threshold Adaptive Encoding (TAE) consistently outperforms Moving Window (MW) and Step Forward (SF) methods across multiple evaluation criteria. TAE demonstrates superior signal reconstruction quality, achieving the lowest reconstruction errors across frequency bands and sound classes, while simultaneously maintaining the lowest spike firing rates, indicating enhanced energy efficiency. These findings establish TAE as the most suitable spike encoding approach for environmental sound applications where both accuracy and energy consumption are critical considerations. The superior performance of TAE aligns with previous findings on adaptive thresholding approaches in neuromorphic signal processing [18, 6], extending their applicability from controlled environments to the more challenging domain of environmental acoustics.
The classification results further validate TAE’s effectiveness, achieving the best performance on two of three evaluated datasets. While all spike encoders perform below traditional baselines, this performance gap is expected given the exploratory nature of applying SNNs to environmental sound classification. This gap highlights opportunities for future architectural improvements and corroborates findings from recent neuromorphic audio surveys [20] that identify the need for specialized SNN architectures tailored to audio processing tasks. The performance gap is also consistent with observations in other domains where SNNs require careful co-design of encoding and network architecture to achieve competitive performance [2, 1].
Complementing this, SATRN (Spiking Audio Tagging Robust Network) shows that attention‑equipped SNN architectures can approach CNN‑like accuracy with efficiency on benchmarks including UrbanSound8K, indicating a clear path to close this gap [21]. Accordingly, a promising direction is to pair a encoder like TAE with attention‑based SNNs like SATRN to jointly optimize encoding fidelity and model capacity for noisy, overlapping scenes and resource‑constrained edge deployments [21, 20].

### v4

@@ 1 Introduction
Spiking Neural Networks (SNNs) have emerged as a promising paradigm in neuromorphic computing due to their energy efficiency and biologically inspired processing mechanisms [Malcolm2023]. Unlike traditional artificial neural networks (ANNs), which rely on continuous-valued activations [Eshraghian2023], SNNs process information through discrete spike events, mimicking the behavior of biological neurons [Zhang2022]. To harness SNNs for real-world applications, sensor data need to be represented as spike trains. While neuromorphic sensors are specifically designed to generate spike-based outputs directly [Lichsteiner:2008:JSSC], conventional sensors typically require an additional encoding step. In this context, various spike encoding methods have been proposed, each with distinct characteristics influencing the efficiency and accuracy of SNN-based models [Auge2021]. The selection of an appropriate encoding scheme is crucial, as it directly affects the network’s capability to extract meaningful temporal patterns from input signals [Petro2020].
Although spike encoding techniques have been extensively studied in the context of speech processing [Stewart2023], their application to environmental sound processing remains underexplored. Environmental sounds, such as traffic noise, alarms, and natural ambient sounds, exhibit complex spectral-temporal structures, often containing overlapped acoustic events and substantial background noise [Springer2024]. Unlike speech, which follows structured linguistic patterns, environmental sounds are highly unstructured and vary significantly in duration and frequency content. These unique challenges require a deeper investigation into the suitability of different spike encoding schemes for environmental sound processing.
State-of-the-art environmental sound classification methods predominantly employ convolutional neural networks (CNNs) or other machine learning models applied to spectrogram representations, which effectively capture the inherent time-frequency characteristics of audio signals [guzhov2021esresnet, kong2020panns, singh2024atgnn, pellegrini2023adapting]. However, to the best of our knowledge, no state-of-the-art solution has yet encoded environmental sound datasets using spike-based methods for SNNs. This gap presents a promising opportunity to investigate neuromorphic-inspired approaches that may offer good temporal resolution and enhanced energy efficiency.
@@ 2.1 Datasets
ESC-10: A subset of ESC-50 containing 400 recordings, divided into 10 classes with 40 clips per class. Each clip has a duration of 5 s. The dataset provides a balanced representation of environmental events, including dog barking, rain, sea waves, crying baby, clock ticking, sneezing, helicopter, chainsaw, crowing rooster, and fire crackling. We used all original 5 s recordings without modification [Piczak2015].
UrbanSound8K: The original dataset contains 8,732 recordings distributed across 10 classes (air conditioner, car horn, children playing, dog bark, drilling, engine idling, gun shot, jackhammer, siren, and street music), each clip lasting up to 4 s. To standardize audio length and ensure a fair comparison in the classification analysis, we retained only clips of exactly 4 s. Having the same lengths avoids padding or other manipulations that could bias the results. We also maintained class balance across folds. However, due to an insufficient number of valid 4 s samples, the classes drilling and car horn were removed, resulting in a filtered subset containing 8 classes and 3,768 clips [salamon2014urbansound].
TAU-3Class: Derived from the DCASE 2020 Acoustic Scene Classification task, we used recordings from device A grouped into three categories: indoor, outdoor, and transportation. Each original recording is 10 s. To match lengths with the other datasets, we extracted the central 5 s segment, ensuring retention of the most relevant acoustic information. We subsampled the official development split to balance the classes and simplify the analysis, yielding 1,500 training and 300 test clips (1,800 total) [Mesaros2018].
@@ 2.2 Signal Reconstruction and Multi-band Analysis
The Mel-spectrogram was chosen for its computational efficiency and demonstrated effectiveness [Stewart2023]. While prior work has employed Gammatone or Butterworth filters to compute cochleagrams [ElFerdaoussi2023, Forno2022], our tests indicated higher computational costs for high-frequency content; the Mel-spectrogram provided a more efficient solution for environmental sound analysis.
@@
Reconstruction quality was assessed using two metrics computed per frequency band—ERRdB (error in decibels) and SNR (signal-to-noise ratio). In addition, ERRdB was computed per sound class for finer-grained analysis. Signal reconstruction was performed using the decoding methods described in [Petro2020].
@@ 2.3 Spike Encoding Methods
Spiking neural networks (SNNs) require transforming continuous signals into discrete spike trains, a process known as spike encoding. The choice of encoding method strongly influences downstream performance. In this study, three techniques—Moving Window (MW), Step Forward (SF), and Threshold Adaptive Encoding (TAE)—were selected based on demonstrated effectiveness in prior work [Forno2022, Petro2020, Wang2024], yet remain underexplored for environmental sound classification, motivating their evaluation here.
@@
Moving Window (MW): Partitions the signal into overlapping windows and emits spikes based on intra-window signal variation, capturing local temporal patterns [Forno2022].
Step Forward (SF): Generates spikes when the difference between consecutive samples exceeds a predefined threshold, preserving salient changes while suppressing minor fluctuations [Forno2022].
Threshold Adaptive Encoding (TAE): Adapts the spike-generation threshold to signal characteristics, improving robustness to amplitude variability and sensitivity to significant changes [Wang2024].
@@ 2.4 SNN Classification
To evaluate classification accuracy, we implemented a spiking neural network (SNN) with four fully connected layers and Leaky Integrate-and-Fire (LIF) neurons using snntorch (v0.9.1) [Eshraghian2023]. The architecture comprises an input layer matching the Mel-spectrogram channels, three hidden layers with 128 neurons each, and an output layer matching the number of classes. Training was performed with a batch size of 32, a learning rate of 0.01, and 100 epochs.
@@ 3.3 Classification
  • Note: Baseline references values taken from: ESC-10 [Piczak2015]; UrbanSound8K [salamon2014urbansound]; TAU-3Class (Task B) [heittola2020task1b].
Our results highlight TAE as the most effective spike encoding method for both reconstruction and classification of environmental sounds, supporting its robust generalization and efficiency. This corroborates prior studies on adaptive thresholding for sequential data [Wang2024] and suggests future work should explore further architectural improvements for challenging sound recognition tasks.
@@ 4 Discussion
Our multi-faceted analysis reveals that Threshold Adaptive Encoding (TAE) consistently outperforms Moving Window (MW) and Step Forward (SF) methods across multiple evaluation criteria. TAE demonstrates superior signal reconstruction quality, achieving the lowest reconstruction errors across frequency bands and sound classes, while simultaneously maintaining the lowest spike firing rates, indicating enhanced energy efficiency. These findings establish TAE as the most suitable spike encoding approach for environmental sound applications where both accuracy and energy consumption are critical considerations. The superior performance of TAE aligns with previous findings on adaptive thresholding approaches in neuromorphic signal processing [Wang2024, Petro2020], extending their applicability from controlled environments to the more challenging domain of environmental acoustics.
The classification results further validate TAE’s effectiveness, achieving the best performance on two of three evaluated datasets. While all spike encoders perform below traditional baselines, this performance gap is expected given the exploratory nature of applying SNNs to environmental sound classification. This gap highlights opportunities for future architectural improvements and corroborates findings from recent neuromorphic audio surveys [Baek2024] that identify the need for specialized SNN architectures tailored to audio processing tasks. The performance gap is also consistent with observations in other domains where SNNs require careful co-design of encoding and network architecture to achieve competitive performance [Eshraghian2023, Malcolm2023].
Complementing this, SATRN (Spiking Audio Tagging Robust Network) shows that attention‑equipped SNN architectures can approach CNN‑like accuracy with efficiency on benchmarks including UrbanSound8K, indicating a clear path to close this gap [SATRN2025]. Accordingly, a promising direction is to pair a encoder like TAE with attention‑based SNNs like SATRN to jointly optimize encoding fidelity and model capacity for noisy, overlapping scenes and resource‑constrained edge deployments [SATRN2025, Baek2024].